OpenStudy (anonymous):

Which table of values goes with the equation y = x^ 2 - 2x + 3? What is the equation of a line that passes through points (0, 4) and (-4, -8)? y= 2x+ 4 y= 2x- 4 y= 3x+ 4 y= 3x- 4

OpenStudy (anonymous):

@Hero @KeganMoore @waterineyes @ganeshie8 @Abarnett help?

OpenStudy (anonymous):

someone please??

OpenStudy (anonymous):

this is two questions, right?

OpenStudy (anonymous):

yes u can answer 1 only

OpenStudy (anonymous):

i can't answer 1 no one can unless you have a choice of table of values

OpenStudy (anonymous):

ok....let me try to attach it!

OpenStudy (anonymous):

i can answer this one What is the equation of a line that passes through points (0, 4) and (-4, -8)

OpenStudy (anonymous):

first find the slope, which is \[\frac{-8-4}{-4-0}=\frac{-12}{-4}=3\]

OpenStudy (anonymous):

and the 2nd one?

OpenStudy (anonymous):

then since \((0,4)\) in on the graph, you know the \(y\) intercept is 4, therefore the equation is \[y=3x+4\]

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

yw